The capacitors of capacitance 6f each are available.  The minimum and maximum capacitance,  which may be obtained

Grade:12th pass

3 Answers

Saurabh Koranglekar
4 years ago
Dear student

Considering capacitors in parallel c = c1+ c2...
C = infinite

Consider series combination
C = (1/6)/(1-(1/6)) = 1/5 farad

Regards
Vikas TU
4 years ago
Dear student .
There is correction in above ans.
If They al are connected in parallel the Cmax = C1+ C2 +.......
It will be = infinite Ceq maximum 
If they are connected in series 
1/Ceq = 1/C1 + 1/C2 + 1/C3 + ….....
1/Ceq = 1/6 + 1/6 + 1/6 …................
This will also tend to infinity , 
But if you connect  2 capacitor in series 
Ceq  minimum = (1/6*1/6)/ (1/6+1/6) = 1/12 Farad 
dushyanta kashyap
4 years ago
For the maximum c1 and c2 , plzz tell me why it is infinite and not equal to c1 + c2 = 12?????????????plzz tell